# What Do Privacy Advertisements Communicate to Consumers?

### v1

@@ 4.4.1. Perceived Actionability of Privacy and Security Advice
Prior to the activity, participants were asked to rank their confidence in being able to perform the steps or go to the setting shown in the campaign on a 5-point scale, with 1 being not at all confident and 5 being extremely confident (Q.18). Apple participants rated themselves as a 4 or 5 in confidence, with most Samsung participants rating themselves at 5. Most Twitter participants rated their confidence at 3. Participants were also asked to rate their motivation to perform the steps shown in the ads or game on a 5-point scale, with 1 being not at all motivated, to 5 extremely motivated (Q.17). Most Apple participants ranked themselves at 3 or 5, and majority of the Samsung participants ranked themselves at 4 or 5. Half of the Twitter participants ranked their motivation at 1. These results could suggest that a negative interaction with the privacy campaign could diminish the participants’ confidence and motivation in enacting privacy protections as recommended by the campaigns.
@@ 4.4.2. Actual Actionability of Privacy and Security Advice
To test actual actionability of privacy and security advice in the campaigns, participants were asked to direct the interviewer to a particular privacy setting or feature that was shown in the campaign (Section 6 & Section 7 of the interview). We asked participants in the Apple treatment group to turn on the “Ask App Not to Track” and “Protect Mail Activity” features, shown in Figure 6. We asked participants in the Samusung group to change location permissions for the Weather App and change microphone permissions for the Zoom app, shown in Figure 7. Finally, we asked Twitter participants to go to the “Ad Preference” setting, shown in Figure 8.
All Apple participants were able to find the “Ask App Not to Track” setting eventually. However, only two participants (A5 and A6, both users) were able to successfully complete the task without help. During this process, A3 requested help twice in order to locate the setting, and A4 used Google to find Apple’s online documentation. In addition, both users and non-users (A1, A2 and A3) did not realize that they had completed the task. The “Ask App Not to Track” feature was presented in the ad as a pop-up that is prompted when opening a newly installed app for the first time. However, it did not show how someone who had already used an app would find the setting, which would have been more consistent with the scenario shown in the video since the protagonist had already installed the apps that were tracking her. Within Apple settings, the setting can be found under Settings $\rightarrow$ Privacy and Security $\rightarrow$ Tracking. This discrepancy was noticed by A5 (a user), who knew she had found the correct setting, adding that “I didn’t know that there was a settings page to manage it. But it does make sense.”
@@
Only one Apple participant (A1) was able to confidently locate the “Protect Mail Activity” setting without help. One participant (A3) did not continue due to having used up their total allocated activity time (3 minutes for the two activities). All remaining participants were able to perform the activity, but either required help or were unsure they had completed the task. During this activity, two participants, a non-user (A4) and a user (A6), both consulted Google to find Apple’s documentation for this feature, with A6 also bringing out their own device to navigate in a more familiar setting. One participant (A5, user) was also unsure whether she had reached the correct setting, and was quickly reassured when the feature had the exact same name as shown in the ad.
All Samsung participants were able to change the location permission for the Weather App. However, only S5 and S6 were able to complete the activity without help. S1 was unsure that she had completed the activity. The device that was used for the Samsung group was a Galaxy A13, and after the desktop was returned to factory settings, the main setting icon is not immediately available on the home screen. The setting icon becomes visible either when one swipes down on the screen, revealing a small gear on the top right corner, or when one swipes up revealing the master setting icon. This default configuration stumped S2, S3, and S5. S3 was a self-described “loyal Samsung customer” who had purchased at least three flagship Galaxy phones in the past. S2 requested help, and S3 used Google twice in an attempt to locate Samsung’s master setting icon. While the setting icon was not available on the home screen, the Google search bar was. S5 directed the interviewer to type “Setting” into the Google search bar and the OS displayed a “From Your Apps: Setting” icon at the bottom of the search, allowing S5 to get into the main setting and progress from there. After the activity, S3 commented that “if I had never seen the Samsung app ad before I would have gone to the right place first. Because I saw the Samsung Ad, I wanted to go to the Samsung privacy first to see if it was in there and then backed out back to the other one.” Both S2 and S3 were unable to continue through to the next activity due to having used up their total allocated activity time.
@@
Finally, for the Twitter group, all participants were able to locate the “Ad Preference Setting,” with T3 and T4 doing so confidently and without help. However, there were minimal connections between participants’ ease of being able to carry out the privacy advice, and what they learned playing the game. T3, T5 and T6 all credited Twitter’s intuitive design, with T5 stating that “That was actually pretty easy. I didn’t know that that was a setting but I feel like where I found it, or where we found it was pretty straightforward.”
@@ 4.4.3. Motivation to Switch or Continue Using the Products
Eight of the 18 study participants who are already users of the device or service said the campaigns motivated them to continue or increase their current usage. A1 said the campaign had no impact on her choice of smartphones, as her decision is based on “other factors.” She said she would stay with Apple because: “I’m not really interested in switching to another brand because it would mess everything up. I’ve got my iCloud calendars and … all that would have to be redone. It would be a huge pain, so I would have to really be motivated by something to try to switch from Apple.” Motivations for continued or increased usage include already being a user and familiarity, like S4 who explained he is motivated to continue using Samsung smartphones because he is familiar with it. Other participants who reported that the campaign motivated them to continue or increase usage referred to feeling reassured from the privacy focused message the company is conveying, with A6 explaining that “[by] bank[ing] their brand on being the data privacy brand … even if [Apple is] not selling my data today, they’re making a public commitment to be the data privacy brand.”
@@
The campaigns only motivated two non-user study participants to consider switching (S2, WD6 for WhatsApp). When asked what motivated them to consider switching, S2 stated “It seemed like it was a simpler process based on the ad the way the phone was set up.” WD6 said the campaign “has me thinking about potentially using that to communicate with some of the people I currently talk to on Snapchat.” In addition, the WhatsApp and DuckDuckGo campaigns inspired a few non-users (WD3, WD6 regarding DuckDuckGo, and WD5 regarding WhatsApp) to look into what these services have to offer. WD3 explained that the advertisement “encouraged me to look into what [DuckDuckGo] is, but not to switch, just to explore what it is.”
For non-users who are not interested in switching, they offered similar reasoning to users, with S5 saying that she would not switch due to “exit cost for me as an iOS user to switch to an Android operating system, like moving all the contacts, or like photos, or like even memos … with the current iOS system, I can already turn on all the features that I am aware of, and it’s just easier for me as a user to keep using my current operating system.” Again, data synchronization and familiarity with their current device and app were credited as the main reasons not to switch. S5 and S6 highlighted the fact that they have similar features in their current smartphones, and the ad did not motivate them to incur the costs of switching from their current smartphone.
@@ 5. Limitations
During the activity phase of the interview, we observed some signs of Hawthorne effect (McCarney et al., 2007), with one Twitter participant stating that they chose not to click on the Twitter setting button as they thought we would like them to focus on the game. Similar observations were also seen in other groups, where non-users often opted not to ask for help even when they were stuck. While we gave reassurance throughout the interview by asking participants to act as they normally would, participants may not have acted so. After the first few participants we began listing examples of “normal” actions, such as using Google or looking on one’s own device, which somewhat alleviated this problem.
@@ 6. Discussion
Increased privacy awareness leads to positive impressions. Our findings suggest that awareness of privacy features contributes to positive perceptions of a company or its products. Campaigns that increased awareness of privacy features improved some participants’ impressions of the associated companies, although they had no impact on the impressions of other participants. In our study, being privacy-focused was credited as a reason for a positive impression of a brand both before and after exposure to the campaigns. This shows that building a privacy focused brand could help building positive impressions. Being suspicious of a company’s privacy practice post exposure to the campaigns could also lead to negative impression, such as T5’s takeaway from DataDash being that Twitter would sell user data.
@@
Avoid invoking skepticism. Campaigns should strive to avoid raising skepticism from viewers and avoid giving conflicting messages. Participants who voiced skepticism towards a company usually expressed concerns about the organization not being able to live up to claims advertised in the campaign, such as questioning whether DuckDuckGo could make money without tracking browsing history. Skepticism towards an organization seemed to be amplified if the participant had prior knowledge about the organization such as exposure to negative news (e.g. data breach), which may conflict with the message that is conveyed in the campaign.
@@
Be careful with visual metaphors. In designing privacy awareness advertisements using visual metaphors, companies need to ensure that viewer comprehension of the metaphor will not arouse negative connotations. One example is Apple’s auctioneer being perceived as representing smartphone companies and possibly Apple itself. This could become a double edge sword that causes viewers to question the company’s intent as they are able to associate the negative connotation as portrayed in the campaign with the organization that issued the campaign. Another more obvious example is when DataDash presented Twitter as a company that the user needed to protect themselves from, which likely goes against any intent Twitter had with launching DataDash. Prior work in the advertising literature cautions advertisers to be careful when selecting visual metaphors, and our results support this (Mohanty and Ratneshwar, 2015, 2016). Creating positive imagery becomes particularly important when viewers are likely to be exposed to the campaign more than once. For those who were previously exposed, continually reinforcing a message can be effective in building a privacy-focused brand image, or allow viewers to pick up on details that may not match the campaign’s message.
@@
Effective privacy ads are short and to the point. We found that the most effective privacy awareness ads were those that were short, focusing on one privacy feature, using a metaphor that is easy to comprehend. This supports Bai et al.’s findings, where the concept of encryption was best communicated when using simple wording with little technical detail (Bai et al., 2020). In our study, we see that participants often walked away with impressions of one main feature. For example, all Apple participants noticed the “Ask App Not to Track” feature that was shown first, while only two noticed the “Protect Mail Activity” feature shown later. Similarly, most participants discussed the location permission control feature in the Samsung ad that was shown first, while only few noticed the camera and microphone permission control features shown later. WhatsApp and DuckDuckGo campaigns were both short and addressed a single feature. Participants seemed to focus their attention on this feature (Encryption for WhatsApp and Private Browsing for DuckDuckGo), and had a good understanding of the campaign messages. Twitter’s DataDash is on the opposite end of the spectrum, where too many abstract elements clouded the campaign’s messages.
@@
DataDash was not effective privacy education. Our results suggests that DataDash was not an effective privacy education game, mainly due to lack of connection between the game and Twitter’s privacy features and setting. In order to use games to educate players about privacy and security concepts, the game should be fun, relatable, and interactive (Sheng et al., 2007; Maqsood et al., 2018). While some of our participants found DataDash fun, the game content might have not been relatable enough to Twitter’s privacy features. Barnard-Wills and Ashenden advised against reducing privacy education games to a “graphical skin overlaid on [an] existing game”, with little connection between the gaming mechanisms and online privacy systems (Barnard-Wills and Ashenden, 2015). Our results support their view, suggesting that DataDash likely fell under this category, with little connection between the game mechanism and Twitter’s privacy features.
@@
Privacy marketing may not be the most effective way to educate users about how to use privacy features. We found that privacy ads and marketing campaigns may not be the best way to show viewers how to use privacy tools and features. The activity segment of the interview highlights the importance of having demonstrations in the ad match what a user will experience. This may be difficult when user experiences may vary. On the other hand, if the campaign makes users aware of a feature and conveys that it is easy to use, users may be motivated to figure out how to use that feature on their own.
@@
To make privacy advice shown more actionable, what is shown in the campaign should remain closely aligned with the actual product or service. From the Apple activity, we see that the difference between the “Ask App Not to Track” pop-up users receive after installing an app and the “Allow Apps to Request to Track” setting can lead to confusion should viewers decide to explore the settings as shown. Given participants’ ease of finding the settings advertised relied on their familiarity with the OS and whether the settings interface was intuitive, it highlights the importance of alignment between the feature advertised and the actual feature.
@@
Privacy campaigns may be most effective for products that don’t incur switching costs and for building positive brand associations over time. Privacy campaigns may not motivate non-users to switch, but could motivate some non-users to explore the product, and could motivate users to continue usage. Participants said they were not motivated to switch products and pointed out that switching costs can be high, especially for smartphones. However, if people associate a particular brand with privacy, it may sway them to switch brands the next time they need to replace their device. In addition, for products with lower switching costs that can be used on a trial basis without giving up an existing service (e.g. messaging apps), privacy campaigns may have a more immediate impact, encouraging viewers to try a service.
@@ 8. Acknowledgment
Eman Alashwali acknowledges the financial support of the Ibn Rushd Program at King Abdullah University of Science and Technology (KAUST). This work was funded in part by Innovators Network Foundation. The authors thank the participants for their time and valuable insights.
@@ A.1. Screening Survey
Answers: $\circ$ “Some high school, no degree” $\circ$ “High school graduate” $\circ$ “Some college, no degree” $\circ$ “Associate’s degree” $\circ$ “Bachelor’s degree” $\circ$ “Graduate degree (Masters, Doctorate, etc.” $\circ$ “Other (Please specify)”
@@
[If Q.5 answer is “No”, the participant skips the rest of this section and moves to Section 3]
@@
[If Q.13 answer is “No”, the participant skips the rest of this section, and moves to Section 5]
@@
[If Q.14 answer does not include “Twitter”, the participant skips the rest of this section, and moves to Section 5]
@@
[If Q.16 answer is “No”, the participant skips the rest of this section, and moves to Section 6]
@@
[If Q.17 answer does not include “WhatsApp”, the participant skips Q.18 and moves to Section 6]
@@
[If Q.19 answer is “No”, the participant skips the rest of this section, and the survey is completed]
@@
[If Q.20 answer does not include “DuckDuckGo”, the participant skips Q.21 and moves to Q.22]
@@
[If Q.22 answer does not include “DuckDuckGo”, the participant skips to Q.23, and the survey is completed]

### v2

@@ 4.4.1. Perceived Actionability of Privacy and Security Advice
Before the activity, participants were asked to rank their confidence in being able to perform the steps or go to the setting shown in the campaign on a 5-point scale, with 1 being “Not at all confident” and 5 being “Fully confident” (Q.18). Apple participants rated themselves as a 4 or 5 in confidence, with most Samsung participants rating themselves at 5. Most Twitter participants rated their confidence at 3. Participants were also asked to rate their motivation to perform the steps shown in the ads or game on a 5-point scale, with 1 being “Not at all motivated” to 5 “Extremely motivated” (Q.17). Most Apple participants ranked themselves at 3 or 5, and the majority of the Samsung participants ranked themselves at 4 or 5. Half of the Twitter participants ranked their motivation at 1. These results suggest that a negative interaction with the privacy campaign could diminish the participants’ confidence and motivation to enact privacy protections as recommended by the campaigns.
@@ 4.4.2. Actual Actionability of Privacy and Security Advice
To test the actual actionability of privacy and security advice in the campaigns, participants were asked to direct the interviewer to a particular privacy setting or feature that was shown in the campaign (Section 6 & Section 7 of the interview). We asked participants in the Apple treatment group to turn on the “Ask App Not to Track” and “Protect Mail Activity” features, shown in Figure 6. We asked participants in the Samsusung group to change location permissions for the Weather App and change microphone permissions for the Zoom app, as shown in Figure 7. Finally, we asked Twitter participants to go to the “Ad Preference” setting, as shown in Figure 8.
All Apple participants were able to find the “Ask App Not to Track” setting eventually. However, only two participants (A5, A6, both users) were able to successfully complete the task without help. During this process, A3 requested help twice to locate the setting, and A4 used Google to find Apple’s online documentation. In addition, both users and non-users (A1, A2, A3) did not realize that they had completed the task. The “Ask App Not to Track” feature was presented in the ad as a pop-up that is prompted when opening a newly installed app for the first time. However, it did not show how someone who had already used an app would find the setting, which would have been more consistent with the scenario shown in the video since the protagonist had already installed the apps that were tracking her. Within Apple settings, the setting can be found under Settings $\rightarrow$ Privacy and Security $\rightarrow$ Tracking. This discrepancy was noticed by A5 (a user), who knew she had found the correct setting, adding: “I didn’t know that there was a settings page to manage it. But it does make sense.”
@@
Only one Apple participant (A1) was able to confidently locate the “Protect Mail Activity” setting without help. One participant (A3) did not continue due to having used up their total allocated activity time (3 minutes for the two activities). All remaining participants were able to perform the activity, but either required help or were unsure they had completed the task. During this activity, two participants, a non-user (A4) and a user (A6), both consulted Google to find Apple’s documentation for this feature, with A6 also bringing out their own devices to navigate in a more familiar setting. One participant (A5, user) was also unsure whether she had reached the correct setting, and was quickly reassured when the feature had the same name as shown in the ad.
All Samsung participants were able to change the location permission for the Weather App. However, only S5 and S6 were able to complete the activity without help. S1 was unsure that she had completed the activity. The device that was used for the Samsung group was a Galaxy A13, and after the desktop was returned to factory settings, the main setting icon was not immediately available on the home screen. The setting icon becomes visible either when one swipes down on the screen, revealing a small gear on the top right corner, or when one swipes up revealing the master setting icon. This default configuration stumped S2, S3, and S5. S3 was a self-described “loyal Samsung customer” who had purchased at least three flagship Galaxy phones in the past. S2 requested help, and S3 used Google twice in an attempt to locate Samsung’s master setting icon. While the setting icon was not available on the home screen, the Google search bar was. S5 directed the interviewer to type “Setting” into the Google search bar and the OS displayed a “From Your Apps: Setting” icon at the bottom of the search, allowing S5 to get into the main setting and progress from there. After the activity, S3 commented: “if I had never seen the Samsung app ad before I would have gone to the right place first. Because I saw the Samsung ad, I wanted to go to the Samsung privacy first to see if it was in there and then backed out back to the other one.” Both S2 and S3 were unable to continue through to the next activity due to having used up their total allocated activity time.
@@
Finally, for the Twitter group, all participants were able to locate the “Ad Preference Setting,” with T3 and T4 doing so confidently and without help. However, there were minimal connections between participants’ ease of being able to carry out the privacy advice and what they learned playing the game. T3, T5, and T6 all credited Twitter’s intuitive design, with T5 stating: “That was actually pretty easy. I didn’t know that that was a setting but I feel like where I found it, or where we found it was pretty straightforward.”
@@ 4.4.3. Motivation to Switch or Continue Using the Products
Eight of the 18 participants who are already users of the device or service said the campaigns motivated them to continue or increase their current usage. A1 said the campaign had no impact on her choice of smartphones, as her decision is based on “other factors.” She said she would stay with Apple because: “I’m not really interested in switching to another brand because it would mess everything up. I’ve got my iCloud calendars and … all that would have to be redone. It would be a huge pain, so I would have to really be motivated by something to try to switch from Apple.” Motivations for continued or increased usage include already being a user and familiarity. For example, S4 explained he is motivated to continue using Samsung smartphones because he is familiar with them. Other participants who reported that the campaign motivated them to continue or increase usage referred to feeling reassured by the privacy-focused message the company is conveying, with A6 explaining that “[by] bank[ing] their brand on being the data privacy brand … even if [Apple is] not selling my data today, they’re making a public commitment to be the data privacy brand.”
@@
The campaigns only motivated two non-user study participants to consider switching (S2, WD6 for WhatsApp). When asked what motivated them to consider switching, S2 stated: “It seemed like it was a simpler process based on the ad the way the phone was set up.” WD6 said the campaign “has me thinking about potentially using that to communicate with some of the people I currently talk to on Snapchat.” In addition, the WhatsApp and DuckDuckGo campaigns inspired a few non-users (WD3, WD6 regarding DuckDuckGo, WD5 regarding WhatsApp) to look into what these services have to offer. WD3 explained that the advertisement “encouraged me to look into what [DuckDuckGo] is, but not to switch, just to explore what it is.”
Non-users who are not interested in switching offered similar reasoning to users, with S5 saying that she would not switch due to “exit cost for me as an iOS user to switch to an Android operating system, like moving all the contacts, or like photos, or like even memos … with the current iOS system, I can already turn on all the features that I am aware of, and it’s just easier for me as a user to keep using my current operating system.” Again, data synchronization and familiarity with their current device and app were credited as the main reasons not to switch. S5 and S6 highlighted the fact that they have similar features in their current smartphones, and the ad did not motivate them to incur the costs of switching from their current smartphone.
@@ 5. Limitations
During the activity phase of the interview, we observed some signs of the Hawthorne effect (McCarney et al., 2007), with one Twitter participant stating that they chose not to click on the Twitter setting button as they thought we would like them to focus on the game. Similar observations were also seen in other groups, where non-users often opted not to ask for help even when they were stuck. While we gave reassurance throughout the interview by asking participants to act as they normally would, participants may not have acted so. After the first few participants, we began listing examples of “normal” actions, such as using Google or looking on one’s own device, which somewhat alleviated this problem.
@@ 6. Discussion
Increased privacy awareness leads to positive impressions. Our findings suggest that awareness of privacy features contributes to positive perceptions of a company or its products. Campaigns that increased awareness of privacy features improved some participants’ impressions of the associated companies, although they had no impact on the impressions of other participants. In our study, being privacy-focused was credited as a reason for a positive impression of a brand both before and after exposure to the campaigns. This shows that building a privacy-focused brand could help build positive impressions. Being suspicious of a company’s privacy practice post-exposure to the campaigns could also lead to a negative impression, such as T5’s takeaway from DataDash being that Twitter would sell user data.
@@
Avoid invoking skepticism. Campaigns should strive to avoid raising skepticism from viewers and avoid giving conflicting messages. Participants who voiced skepticism towards a company usually expressed concerns about the organization not being able to live up to claims advertised in the campaign, such as questioning whether DuckDuckGo could make money without tracking browsing history. Skepticism towards an organization seemed to be amplified if the participant had prior knowledge about the organization, such as exposure to negative news (e.g., data breach), which may conflict with the message that is conveyed in the campaign.
@@
Be careful with visual metaphors. In designing privacy awareness advertisements using visual metaphors, companies need to ensure that viewer comprehension of the metaphor will not arouse negative connotations. One example is Apple’s auctioneer being perceived as representing smartphone companies and possibly Apple itself. This could become a double-edge sword that causes viewers to question the company’s intent as they can associate the negative connotation as portrayed in the campaign with the organization that issued the campaign. Another more obvious example is when DataDash presented Twitter as a company that the user needed to protect themselves from, which likely goes against any intent Twitter had with launching DataDash. Prior work in the advertising literature cautions advertisers to be careful when selecting visual metaphors, and our results support this (Mohanty and Ratneshwar, 2015, 2016). Creating positive imagery becomes particularly important when viewers are likely to be exposed to the campaign more than once. For those who were previously exposed, continually reinforcing a message can be effective in building a privacy-focused brand image or allowing viewers to pick up on details that may not match the campaign’s message.
@@
Effective privacy ads are short and to the point. We found that the most effective privacy awareness ads were those that were short, focusing on one privacy feature, using a metaphor that is easy to comprehend. This is consistent with Bai et al.’s finding that the concept of encryption was best communicated when using simple wording with little technical detail (Bai et al., 2020). In our study, we see that participants often walked away with impressions of one main feature. For example, all Apple participants noticed the “Ask App Not to Track” feature that was shown first, while only two noticed the “Protect Mail Activity” feature shown later. Similarly, most participants discussed the location permission control feature in the Samsung ad that was shown first, while only a few noticed the camera and microphone permission control features shown later. WhatsApp and DuckDuckGo campaigns were both short and addressed a single feature. Participants seemed to focus their attention on this feature (Encryption for WhatsApp and Private Browsing for DuckDuckGo) and had a good understanding of the campaign messages. Twitter’s DataDash is on the opposite end of the spectrum, where too many abstract elements clouded the campaign’s messages.
@@
DataDash was not effective in privacy education. Our results suggest that DataDash was not an effective privacy education game, mainly due to a lack of connection between the game and Twitter’s privacy features and settings. To use games to educate players about privacy and security concepts, the game should be fun, relatable, and interactive (Sheng et al., 2007; Maqsood et al., 2018). While some of our participants found DataDash fun, participants had trouble relating the game content to Twitter’s privacy features. Barnard-Wills and Ashenden advised against reducing privacy education games to a “graphical skin overlaid on [an] existing game,” with little connection between the gaming mechanisms and online privacy systems (Barnard-Wills and Ashenden, 2015). Our results support their view, suggesting that DataDash likely fell under this category, with little connection between the game mechanism and Twitter’s privacy features.
@@
Privacy marketing may not be the most effective way to educate users about how to use privacy features. We found that privacy ads and marketing campaigns may not be the best way to show viewers how to use privacy tools and features. The activity segment of the interview highlights the importance of having demonstrations in the ad to match what a user will experience. This may be difficult when user experiences may vary. On the other hand, if the campaign makes users aware of a feature and conveys that it is easy to use, users may be motivated to figure out how to use that feature on their own.
@@
To make privacy advice more actionable, what is shown in the campaign should remain closely aligned with the actual product or service. From the Apple activity, we see that the difference between the “Ask App Not to Track” pop-up users receive after installing an app and the “Allow Apps to Request to Track” setting can lead to confusion should viewers decide to explore the settings as shown. Given participants’ ease of finding the settings advertised relied on their familiarity with the OS and whether the settings interface was intuitive, it highlights the importance of alignment between the feature advertised and the actual feature.
@@
Privacy campaigns may be most effective for products that don’t incur switching costs and for building positive brand associations over time. Privacy campaigns may not motivate non-users to switch, but could motivate some non-users to explore the product, and could motivate users to continue usage. Participants said they were not motivated to switch products and pointed out that switching costs can be high, especially for smartphones. However, if people associate a particular brand with privacy, it may sway them to switch brands the next time they need to replace their device. In addition, for products with lower switching costs that can be used on a trial basis without giving up an existing service (e.g., messaging apps), privacy campaigns may have a more immediate impact, encouraging viewers to try a service.
@@ 8. Acknowledgment
Eman Alashwali acknowledges the financial support of the Ibn Rushd Program at King Abdullah University of Science and Technology (KAUST). This work was funded in part by the Innovators Network Foundation. The authors thank the participants for their time and valuable insights.
@@ A.1. Screening Survey
Answers: $\circ$ “Some high school, no degree” $\circ$ “High school graduate” $\circ$ “Some college, no degree” $\circ$ “Associate’s degree” $\circ$ “Bachelor’s degree” $\circ$ “Graduate degree (Masters, Doctorate, etc.)” $\circ$ “Other (Please specify)”
@@
[If Q.5 answer is “No,” the participant skips the rest of this section and moves to Section 3]
@@
[If Q.13 answer is “No,” the participant skips the rest of this section, and moves to Section 5]
@@
[If Q.14 answer does not include “Twitter,” the participant skips the rest of this section, and moves to Section 5]
@@
[If Q.16 answer is “No,” the participant skips the rest of this section, and moves to Section 6]
@@
[If Q.17 answer does not include “WhatsApp,” the participant skips Q.18 and moves to Section 6]
@@
[If Q.19 answer is “No,” the participant skips the rest of this section, and the survey is completed]
@@
[If Q.20 answer does not include “DuckDuckGo,” the participant skips Q.21 and moves to Q.22]
@@
[If Q.22 answer does not include “DuckDuckGo,” the participant skips Q.23, and the survey is completed]